OpenStudy (anonymous):

Find the limit of ln (1-x^2)/lnx when x tends to infinity

OpenStudy (shubhamsrg):

undefined ?

OpenStudy (shubhamsrg):

@sauravshakya thats the wrong identity ..

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

Ya..... thats wrong

OpenStudy (shubhamsrg):

well log(-ve) is undefined,,so wont the limit, simply not exist..?

hartnn (hartnn):

u know L'hopital's rule ? @duwduuw

OpenStudy (anonymous):

2!

OpenStudy (anonymous):

Ya.... @hartnn has correct approach

OpenStudy (anonymous):

(not factorial, merely excitement)

OpenStudy (anonymous):

I think its 0

OpenStudy (anonymous):

\[\frac{ \log(1-x )+\log_{}(1+x ) }{ logx }\] evaluate seperately

hartnn (hartnn):

i was thinking exactly that jonask

OpenStudy (anonymous):

Yep i do know the hospitals rule with the derivatives!

OpenStudy (anonymous):

are they difficult graphs?

hartnn (hartnn):

so did u try differentiating numerator and denominator separately ?

OpenStudy (anonymous):

"the hospitals rule" ^.^

OpenStudy (anonymous):

Again got the same.... its 0

OpenStudy (anonymous):

\[\frac{ \infty+\infty }{ \infty }=2\] is there such thing

hartnn (hartnn):

did u get (2x^2)/(1-x^2) ?? and answer = 2

OpenStudy (anonymous):

@Jonask there is no such thing

OpenStudy (shubhamsrg):

log(1-x^2) at x-->infinity = log(-infinity) -->undefined,,i'd simply stick to my ans that it is undefined..

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

One week ago, every time a limit showed up, everyone immediately screamed "l'Hopital's !". Now it seems we work up to it slowly.

OpenStudy (anonymous):

@shubhamsrg is correct

hartnn (hartnn):

wolf says its 2 !!!

OpenStudy (anonymous):

@shubhamsrg is correct

OpenStudy (anonymous):

You're all fired. Except @hartnn . He's promoted to assistant to the regional manager.

OpenStudy (anonymous):

please stop joking @Algebraic!

OpenStudy (anonymous):

graphycally

OpenStudy (anonymous):

use l'Hopital's /thread

hartnn (hartnn):

@duwduuw u differentiated numerator and denominator, right ? what did u get ?

OpenStudy (shubhamsrg):

you sure the question doesnt have ln(x^2 -1) there?? had it been that,,surely ans is 2 then.. otherwise ..you all know my answer! :P

OpenStudy (anonymous):

2x/xln(1+x^2)

OpenStudy (experimentx):

\[ \ln (1 + x) = \ln x + \ln (1/x + 1) \\ \ln(1 - x) = \ln(-x) + \ln (-1/x + 1)\]

OpenStudy (experimentx):

interesting identity http://www.wolframalpha.com/input/?i=limit+x-%3Einf+ln%28-x%29%2Fln%28x%29 L'hopital is much simpler.

OpenStudy (shubhamsrg):

ohh!! :O

hartnn (hartnn):

ln term will not come in differentiating num and denom separately. u get (2x^2)/(1-x^2)

hartnn (hartnn):

even there L'Hopital's is used !

OpenStudy (shubhamsrg):

since log(-1) = i(pi) ,,,its some constant,,but still can we ignore it in front of infinity ?

OpenStudy (experimentx):

you can do that without L'hopital's rule ... you need to make a slight modification of this identity http://www.wolframalpha.com/input/?i=limit+x-%3E0+%28-x%29^x

OpenStudy (shubhamsrg):

aha.. @mukushla will you confirm this please ?

OpenStudy (anonymous):

Oooooo nevermind, i meant to say 2x^2/1+x^2

OpenStudy (anonymous):

I agree hartnn

hartnn (hartnn):

u get 2x^2/(1-x^2) = 2/(1/x^2-1) and when u put x=infinity, u get 2.

OpenStudy (anonymous):

Exactly, so 2 is the limit. +1 for hartnn haha

hartnn (hartnn):

thank u :)

OpenStudy (anonymous):

What am I? Chopped liver?

OpenStudy (anonymous):

daw

OpenStudy (anonymous):

Haha sorry Just saw that you had the answer firstttt